# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 4 Sublinear convergence of stochastic PDR for restricted primal-dual gap
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2}\Big{[}\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|_{\frac{1}{\sigma}}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}\Big{]}$
$+\frac{1}{2}(\Delta^{0}+\frac{1}{\underline{p}}M_{K})+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})$
$-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Here $C$ is produced in Lemma 5, $M_{K}$ is from (4.13), $\Delta^{0}=\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$, $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$, $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{Y}$ randomly. The same rate also holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}(z^{K}_{\operatorname{test}},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]$
$+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))].$ (4.21)
@@
and two diagonal weights $\mathbf{w}_{x}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\sigma}I]$, $\mathbf{w}_{y}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I]$. Therefore, $\|u^{k}-u\|^{2}_{\mathbf{M_{p}}}=\|\mathbf{x}^{k}-\mathbf{x}\|^{2}_{\mathbf{w}_{x}}+\|\mathbf{y}^{k}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}$, in which $\mathbf{w}_{y}\mathbf{P}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I)\mathbf{P},\tfrac{1}{\tau}\mathbf{P}]$. For $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, we focus on the iteration $u^{k}\rightarrow\tilde{u}^{k+1}$. Combining Proposition 1 and (3.2), we obtain
@@
$\begin{array}[]{c}\Delta_{\mathbf{x}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}}^{2}\Big{)},\\ \Delta_{\mathbf{y}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\Big{)}.\\ \end{array}$ (4.24)
@@
$\Delta_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}+\|\mathbf{y}^{k}-\ y\|_{\mathbf{w}_{y}}^{2}$
$+(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}$
$=-\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}+\epsilon_{\mathbf{y}}^{k},$ (4.25)
@@
$\epsilon_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}$
$=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-2\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}\Big{]}.$ (4.26)
@@
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}}^{2}$
$+\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\Big{)}+\epsilon_{\mathbf{y}}^{k}$
$=\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{p}}-\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p}}-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}\Big{)}$
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}+\frac{1}{2}\Big{(}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}}$
$+\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\textbf{}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}\Big{)},$ (4.27)
where the equality is by adding and subtracting $\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$ and the definition of $\epsilon_{\mathbf{y}}^{k}$ in (4.26). In (4.27), dropping $-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}$ and summing over $k=0,\ldots,K$ yields,
$\sum^{K}_{k=0}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}}+\frac{1}{2}\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\mathcal{E}^{K}+\sum_{k=0}^{K}[-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}],$ (4.28)
where $\mathcal{E}^{K}$ is defined as follows
$\frac{1}{2}\sum_{k=0}^{K}\left[(\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}})+(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\right].$
@@
$\mathbb{E}\mathcal{E}^{K}/K$ $=\frac{1}{2K}\mathbb{E}\sum_{k=0}^{K}\Big{[}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\mathbb{E}^{k}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$
$+\mathbb{E}^{k}(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}=0.$ (4.29)
@@
For the sum of inner products in (4.28), introducing $\mathbf{y}^{k+1}_{\dagger}=\mathbf{y}^{k}_{\dagger}-\mathbf{P}\mathbf{v}^{k+1}$ and selecting a $\mathbf{y}^{0}_{\dagger}\in\mathbb{B}_{2}\times\mathbb{B}_{2}$ randomly, we have
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}=\langle\mathbf{y}_{\dagger}^{k}-\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}.$
@@
In (4.28), dropping $-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}}$, dividing both sides by $K$, taking supremum with respect to $z$, and employing (4.14) on the above inner product, we arrive at

### v3

@@ 4 Sublinear convergence of stochastic RPDR for restricted primal-dual gap
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2\rho_{l}}\Big{[}\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\Big{]}+\sup_{z\in\mathbb{B}}\frac{1}{2}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2})$
$+\frac{\rho_{u}^{2}}{2(2-\rho_{u})\rho^{2}_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}}M_{K}$
$+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Here $C$ is produced in Lemma 5, $M_{K}$ is from (4.13), $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$, $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{B}_{2}$ randomly. The same rate also holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}(z^{K}_{\operatorname{test}},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]$
$+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))].$ (4.21)
@@
and the following diagonal weights $\mathbf{w}_{x}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\sigma}I]$, $\mathbf{w}_{y}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I]$, $\mathbf{R}_{x}^{k}:=\operatorname{Diag}[\tfrac{1}{\rho_{k}}I,\tfrac{1}{\rho_{k,x}}I]$, $\mathbf{R}_{y}^{k}:=\operatorname{Diag}[\tfrac{1}{\rho_{k}}I,\tfrac{1}{\rho_{k,y}}I]$, $\mathbf{r}_{x}^{k}:=\operatorname{Diag}[\frac{2-\rho_{k}}{\rho_{k}^{2}}I,\frac{2-\rho_{k,x}}{\rho_{k,x}^{2}}I]$, and $\mathbf{r}_{y}^{k}:=\operatorname{Diag}[\frac{2-\rho_{k}}{\rho_{k}^{2}}I,\frac{2-\rho_{k,y}}{\rho_{k,y}^{2}}I]$. Therefore, $\|u^{k}-u\|^{2}_{\mathbf{M_{p}}}=\|\mathbf{x}^{k}-\mathbf{x}\|^{2}_{\mathbf{w}_{x}}+\|\mathbf{y}^{k}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}$, in which $\mathbf{w}_{y}\mathbf{P}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I)\mathbf{P},\tfrac{1}{\tau}\mathbf{P}]$. For $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, we focus on the iteration $u^{k}\rightarrow\tilde{u}^{k+1}$. Combining Proposition 1 and (3.3), we obtain
@@
$\begin{array}[]{c}\Delta_{\mathbf{x}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}\mathbf{r}_{x}^{k}}^{2}\Big{)},\\ \Delta_{\mathbf{y}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}\Big{)}.\\ \end{array}$ (4.24)
@@
$\Delta_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k}-\ y\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}$
$+(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}$
$=-\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\frac{1}{2}\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}+\epsilon_{\mathbf{y}}^{k},$ (4.25)
@@
$\epsilon_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}$
$=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-2\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}\Big{]}.$ (4.26)
@@
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}\mathbf{r}_{x}^{k}}^{2}$
$+\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}\Big{)}+\epsilon_{\mathbf{y}}^{k}$
$=\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}\Big{)}$
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}+\frac{1}{2}\Big{(}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}$
$+\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}\Big{)},$ (4.27)
where the equality is by adding and subtracting $\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}$ and the definition of $\epsilon_{\mathbf{y}}^{k}$ in (4.26). In (4.27), dropping $-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}}$ and summing over $k=0,\ldots,K$ yields,
$\sum^{K}_{k=0}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{K}}}+\frac{1}{2}\|u^{0}-u\|^{2}_{\mathbf{M}_{p,\varrho_{0}}}+\mathcal{E}^{K}+\sum_{k=0}^{K}[-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}],$ (4.28)
where we use $\mathbf{M}_{p,\varrho_{k+1}}\leq\mathbf{M}_{p,\varrho_{k}}$ for all $k=0,\dots,K$. $\mathcal{E}^{K}$ is defined as follows
$\frac{1}{2}\sum_{k=0}^{K}\Big{[}$ $(\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}})$
$+(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}.$
@@
$\mathbb{E}\mathcal{E}^{K}/K_{1}$ $=\frac{1}{2K_{1}}\mathbb{E}\sum_{k=0}^{K}\Big{[}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\mathbb{E}^{k}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}$
$+\mathbb{E}^{k}(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}=0.$ (4.29)
@@
For the sum of inner products in (4.28), introducing $\mathbf{y}^{k+1}_{\dagger}=\mathbf{y}^{k}_{\dagger}-\mathbf{P}\mathbf{v}^{k+1}$ and selecting a $\mathbf{y}^{0}_{\dagger}=(y^{0}_{\dagger},\bar{y}^{0}_{\dagger})\in\mathbb{B}_{2}\times\mathbb{B}_{2}$ randomly, we have
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}=\langle\mathbf{y}_{\dagger}^{k}-\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}.$
@@
In (4.28), dropping $-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}\mathbf{R}^{k}}$, dividing both sides by $K_{1}$, taking supremum with respect to $z$, and employing (4.14) on the above inner product, we arrive at